A triangle drawn on a map has sides of lengths 7 ​cm, 12 ​cm, and 14 cm. The shortest of the corresponding ​real-life distances is 112 km. Find the longest of the ​real-life distances. Round to the nearest tenth as needed.
Answer by jim_thompson5910(35256)


a = shortest distance on paper = 7 cm
b = shortest distance in real life = 112 km

c = longest distance on paper = 14 cm
d = longest distance in real life = x km

The corresponding sides form a proportion.
a/b = c/d
7/112 = 14/x ... plug in the values for a through d (see above)
7*x = 112*14 ... cross multiply
7x = 1568
7x/7 = 1568/7 ... divide both sides by 7
x = 224

Answer: 224 kilometers